I am unable to figure out a formula for the following : Sheila's age is 2 years more than twice Nicole's age. The sum of sheila and Nicole's ages is the same as Steve's. Steve's age is ten years less than Nicole's. What are all three ages ?
Thank you for your help !

Answer by ankor@dixie-net.com(22740)

let s = Sheila, present age
let n = Nicole's age
let v = Steve's age
:
Write an equation for each statement:
" Sheila's age is 2 years more than twice Nicole's age."
s = 2n + 2
" The sum of sheila and Nicole's ages is the same as Steve's."
s + n = v
"Steve's age is ten years less than Nicole's."
v = n-10
Right here, looking at the last two equations we can see this is not
a possible situation, look at this:
v = v so we have
s + n = n - 10
s = -10, a negative age
A bad word problem


What are all three ages ?